Status Report

ISS On-orbit Status 14 Jan 2003

By SpaceRef Editor
January 14, 2003
Filed under , ,
ISS On-orbit Status 14 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.  A relatively light day for the crew, before tomorrow’s spacewalk by CDR Bowersox and FE-2/SO Pettit (no sleep cycle shift).

EVA preparations are going well, and yesterday’s prep dry-runs were completed by all crewmembers with excellent results.  [They confirmed good EMU fit, and the CDR also practiced holding FE-2’s legs in the Airlock (A/L) as Pettit simulated performing the Node ACBM (active common berthing module) cleaning task.]

EVA Objectives:  Prior to the EVA (egress: 7:30am EST), the MT (mobile transporter) will be moved from WS7 (work station 7) at the end of the P1 truss to WS4, allowing sufficient time for crew assist if required (to install MT extension cord).  Planned EVA tasks are:  (1) configure SFU (squib firing unit) on P1 to launch condition; (2) release 10 launch restraint locks (5 zenith, 5 nadir) on the P1 radiator beam; (3) inspect radiator beam and monitor its deployment (with manual intervention if required); (4) reconfigure SFU; (5) clean Node nadir ACBM (active common berthing module); (7) preposition an APFR (articulating portable foot restraint) on the CETA (crew equipment transfer aid) cart, and (8) perform get-ahead tasks as open time permits.  It will be the 50th EVA for ISS assembly, with 25 from Shuttle and 25 from station (16 of these from the Joint Airlock, but it will be only the second with no Orbiter present).

Today, the crew had a final tag-up with ground specialists via S-band to review EVA timeline and procedures.  [The review included late timeline changes as well as contingency cases, e.g., steps to be taken in the event of LOAC (loss of CMG attitude control) during the MT (mobile transporter) translation from WS7 (work station 7) to WS4 and P1 radiator deployment, when the Russian segment thrusters are disabled (in which case MT or radiator activity would immediately be stopped and safed).  Also discussed were onboard procedures required in case of loss of communications with the spacewalkers or MCC-H.]
CDR Ken Bowersox installed Metox (metal oxide) canisters and batteries in the EMUs, one each in suit #3013 (CDR) and suit #3011 (FE-2).  Also, both spacewalkers prepared EVA hardware in the A/L equipment compartment.

FE-1 Nikolai Budarin completed his daily routine inspection of the Russian BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) experiment, which studies plant growth in the zero-G environment.

Budarin also prepared the commentary and photo documentation files of yesterday’s windows inspection in the Service Module (SM) and Docking Compartment (DC-1) for downlink to MCC-M.

Later, Nikolai worked on the SM’s ASU toilet system, replacing the urine receptacle filter insert (F-V).

FE-2/SO Pettit continued his daily monitoring of the ACS (atmosphere control and supply) system CO2 (carbon dioxide) levels in SM and Lab with the CDM (CO2 monitor).

For EVA video coverage, Pettit set up and checked out the VTR (video tape recorder) bypass cable configuration on the CVIU (common video interface unit) to provide additional camera views on two separate Sony V10 monitors located at the Lab RWS (robotic work station).  [Necessary VDS (video distribution system) routing to two VTRs and VBSPs (video baseband signal processors) was performed via S-band by the ground.]

Don Pettit also completed the regular weekly maintenance of the TVIS (treadmill with vibration isolation and stabilization), as usual at the end of his exercise.  [For their exercise on the treadmill, crewmembers are still required to use the SPDs (subject positioning devices) and to limit their running speed to no more than 6 mph.]

The daily routine tasks of SOSh life support systems maintenance and Lab autonomous payload status checkup were conducted by Bowersox, while Pettit prepared the daily IMS inventory delta file for automated export to MCC-M.

The S1 truss radiator beam was rotated by remote control on its TRRJ (thermal radiator rotary joint) to an angle of 20 deg to provide clearance for EVA translation.

After the switch of the Lab CCAA (common cabin air assembly) from starboard to port air conditioner on 1/12 in conjunction with CDRA (CO2 removal assembly) activation, the port heat exchanger has experienced an increase in the frequency of “wet” indications; two indications occurred yesterday.  The Lab temperature set point has now been increased to preclude further “wet” indications, and the A/L CCAA has been activated to assist.  The Lab CCAA set point will be gradually reduced to the crew-requested temperature.

This morning, the crew inspected the failed Vozdukh CO2 scrubber and successfully restarted the system.  At this time, Vozdukh is up and running again.  [Even so, due to a failed timing unit Vozdukh can operate only in manual cycle mode (i.e., manual switching between filter beds) until at least 2/4 when Progress 10P will deliver a spare for the failed timer (which is required for automatic cycle mode).]
MCA (major constituents analyzer) continues failed, but investigation is underway.  Failure signature indicates aging ion pump (which, however, was recently installed).  MCA is “nice to have” during EVA, but not a requirement.

Today, three new U.S. HDDs (hard disk drives) were scheduled to be carried  to Baikonur for loading on Progress 10P.  [The HDDs contain software serving as crew interface for the next station software upgrade next month.  A meeting on the planned s/w upgrade was held at JSC today.]

Water Processor (SRV-K) Update:   It is not clear at this time when the Russian SRV-K water processing system will again be able to process U.S. condensate water after Progress 10P.  [Troubleshooting of the basic problem of the SRV-K, whose separator unit is getting “overwhelmed” with water, has not been successful to date.  Some new spares will fly on 10P, including additional Russian EDV water container bags for storing excess U.S. condensate, but RSC-Energia’s solution to the basic problem may be a new, more capable CFU (condensate feed unit) of the Elektron O2 generator, along with bigger pipes and hoses.  With the 10P SORR (Stage Operations Readiness Review) coming up on 1/16 (Thursday), a decision by RSC-E is required early enough to allow U.S. planning for water up-cargo on Mission ULF-1 in early March.  Meanwhile, both U.S. condensate CWCs (contingency water containers) have been topped off, and at least one overboard dump will probably be required in the next few days (details in work). ]

Microgravity Science Glovebox (MSG) Update:  The re-worked PDC (power distribution controller) is expected to arrive at MSFC/Huntsville this weekend for integrated engineering testing and recertification in the ground unit, until 1/20 or 1/21.  Since the second (spare) PDC has problems that prevent its flying on the Progress, the criticality of successful PDC flight unit testing is increased.  ESA plans to have the flight unit in Moscow by 1/23, if possible.  Since RSC-Energia would have difficulties delaying the 2/2 launch, the PDC and ESEM3 (exchangeable standard electronic module 3) must be in Baikonur by 1/25, when 10P is closed out for irreversible liquid loading ops.  [In case the parts “miss the boat”, next launch opportunity would then be on ULF-1, in which case MSG operations would start so late that the European taxi mission on the next Soyuz (6S) in late April may not be able to get MSG time.]

RED (Resistive Exercise Device) Update:  After the recent (1/10) installation of the backup RED canister set by the crew, it has been decided now to bring only one canister back on ULF-1, instead of the originally planned two.  This leaves three canisters on orbit, and two refurbished units, returned on 11A, will be delivered by ULF-1.  Engineers are still developing an IFM (in-flight maintenance) procedure for the on-orbit cans with the pulley-retraction problem.

Today’s targets for the CEO (crew earth observations program) were Lake Victoria, Kenya (Dynamic Event. As clear as it gets.  Crew was to look just right of track. Views of the rapidly changing coastline is of great interest since disturbance on coasts negatively impacts fishing which supports millions of people), Nairobi, Kenya (nadir and a touch left; ESC [electronic still camera]), Dar es Salaam, Tanzania (nadir pass over Tanzania’s capital city; ESC), Kinshasa, Congo (Zaire) (nadir and a touch left; ESC), Industrialized SE Africa (aerosol buildup.  Looking right of track across the plateau towards the Witwatersrand industrial region, the largest in Sub-Saharan Africa), Caracas, Venezuela (nadir; ESC), Brasilia, Brazil (nadir; ESC), Sabancaya Landslide, Peru (there is interest in those few points of landside hazard on the steepest slopes of the Andes Mts. Lack of vegetation and atmospheric clarity makes remote sensing a very viable method for analysis of surface geology. Suggest a mapping pass at nadir for about 30 sec), Lake Poopo, Bolivia (this high altitude, non-outlet lake is a prime sensor for snow/rain precipitation in a large basin.  The high plateau generally dries out when El Nino conditions are in place. We expect that the lake is smaller now than it was when last photographed about two months ago), Patagonian Glaciers (handheld photos have covered many of the dozens of glacier tongues with detailed views for the first time.  Some still need to be documented: looking to nadir and slightly left and right of rack), Pearl & Hermes reefs (nadir pass.  Detailed views needed for global map), Lisianski reef, Hawaiian chain (nadir pass.  Detailed views needed for global map), Laysan reef, Hawaiian chain (sun glint opportunity for great detail of coastline), Howland Island reef (nadir pass.  Detailed views needed for global map.  This oval island is 1.5 miles long with a continuous coral reef.  The beacon on the island is Earhart Light), and Baker Island (nadir pass.  Detailed views needed for global map.  The island is 2.5 times the size of The Mall in Washington, D.C.).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (16 Amp mode).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is On. TCCS (trace contaminant control subsystem) is operational.  MCA (major constituents analyzer) is failed.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 734; temperature (deg C) — 23.6; ppO2 (mmHg) — 147.6; ppCO2 (mmHg) — 2.2 (suspect).
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 18.8.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 732.44; temperature (deg C) — 23.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 735.02; temperature (deg C) — 20.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 734.92, temperature (deg C) — 22.6; shell heater temp (deg C) — 21.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.1
  • PMA-2:  Shell heater temp (deg C) — 17.6

(n/a = data not available)

Propulsion System (PS):  Total propellant load available (SM + FGB + Progress) — 3689 kg (8133 lb) [as of 1/9/03].

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B both in Dual-angle mode (directed position).
  • SM batteries (as of 3am):  Battery #7 is off line (failed); battery #4 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries (3am):  Batteries #3 is offline (failed); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby; PCU-2 is in Standby.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off, SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-2 is On (primary), EXT-1 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1
  • Attitude — U.S. SIGI-1
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • LVLH TEA (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.1 deg, roll: 0 deg]).
  • Solar Beta Angle: 16.3 (magnitude increasing)

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operating.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.  MT: at WS7, with KA power.  POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:53am EST [= epoch]):

  • Mean altitude — 390.3 km
  • Apogee — 394.0 km
  • Perigee — 386.6 km
  • Period — 92.36 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005468
  • Orbits per 24-hr. day — 15.59
  • Solar Beta Angle — 16.3 deg (magnitude increasing)
  • Altitude loss — 180 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 23703
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.